In Arecent 159 ARecent issue of Itta engineering conferences 4 Conferences lasted a day 89 engineering co 36 lasted 3 days
Answer #1
X (length in days) Frequency (no. of conf) cumulative freq.
2 4

4

3 36 40
4 18 58
5 19 77
6 4 81
7 1 82
8 1 83
9 1 84
total 84

b) Median : mean of (84/2)th and (84/2 +1 )th value

= mean of 42nd and 43rd value

42nd value is 4 and 43rd value is 4

Thus median = 4

Q1= (84/4) =21st value = 3

c) 65th percentile = 84 * (65/100) =54.6 i.e, 55th value= 4

d) 10th percentile = 84 * (10/100) =8.4 i.e, 9th value= 3

e) Boxplot

Вох plot 2 3 4 6 8 5 7

For the middle 50% of the first and the third quartile as middle 50% of the data lies between first and third quartile

Q1= 3

Q3= (3*84)/4 =63rd data= 5

The middle 50% of the conference id from 3 days to 5 days.

Sample mean =fa/n = ( 4*2 + 36*3 + 18*4+........ + 1*8 + 1*9)/ 84

= 3.94 ~ 4 days